OpenStudy (anonymous):

A set of data is normally distributed with a mean of 100 and standard deviation of 15. · What would be the standard score for a score of 85? · What percentage of scores is between 100 and 85? · What would be the percentile rank for a score of 85?

OpenStudy (anonymous):

A set of data is normally distributed with a mean of 100 and standard deviation of 15. What would be the standard score for a score of 85? z(85) = (85-100)/15 = -15/15 = -1 That means 85 is one standard deviation to the left of the mean.

OpenStudy (anonymous):

Now one standard deviation from the mean covers 34% of the population. normal (85,100,100,15) = 0.3413

OpenStudy (anonymous):

And the lat one , the tail to the left of 85 must have 0.5000-0.3413 = 0.1587 of te population. The percentile score must be 15.87% . got it ?

OpenStudy (anonymous):

yes I understand now thanks ashna

OpenStudy (anonymous):

no mention :)